# PRISM: Prompt-Refined In-Context System Modelling for Financial Retrieval

### v1

@@ A.5.1 Cost Analysis
Table 6: Latency, token usage, and cost statistics for different runs used in feasibility analysis.
@@
ID Task Latency Statistics Token Usage Statistics Cost Statistics (USD) Min $Q_{1}$ Median $Q_{3}$ Max $\sum_{prompt}$ $\sum_{completion}$ $\mu_{prompt}$ $\mu_{completion}$ $\sum_{prompt}$ $\sum_{completion}$ $\sum_{combined}$ 2 Document 2 3 3 4 100 150,541 61,904 752.71 309.52 0.04 0.12 0.16 2 Chunk 4 12 32 42.5 4190 15,010,574 519,830 75,052.87 2,599.15 3.75 1.04 4.79 2 Combined 2 3 7 32.25 4190 15,161,115 581,734 37,902.79 1,454.34 3.79 1.16 4.95 9 Document 2 4 4 5 17 241,941 68,752 1,209.71 343.76 0.06 0.14 0.20 9 Chunk 6 13 32 47 99 14,963,931 490,556 74,819.66 2,452.78 3.74 0.98 4.72 9 Combined 2 4 8 32 99 15,205,872 559,308 38,014.68 1,398.27 3.80 1.12 4.92 12 Document 5 12 17 25 83 327,741 297,498 1,638.71 1,487.49 0.08 0.59 0.68 12 Chunk 9 61.75 135.5 194 665 15,214,202 2,271,311 76,071.01 11,356.56 3.80 4.54 8.35 12 Combined 5 17 37.5 135.25 665 15,541,943 2,568,809 38,854.86 6,422.02 3.89 5.14 9.02 13 Document 2 4 5 6 15 426,941 84,369 2,134.71 421.85 0.11 0.17 0.28 13 Chunk 34 92 105 117 171 19,200,124 1,734,012 96,000.62 8,670.06 4.80 3.47 8.27 13 Combined 2 5 24.5 105 171 19,627,065 1,818,381 49,067.66 4,545.95 4.91 3.64 8.54 15 Document 2 5 7 10 25 426,941 180,699 2,134.71 903.50 0.53 1.81 2.34 15 Chunk 23 132.75 156 179 240 19,170,218 3,159,981 95,851.09 15,799.91 23.96 31.60 55.56 15 Combined 2 7 24 156 240 19,597,159 3,340,680 48,992.90 8,351.70 24.50 33.41 57.90 17 Document 3 5 6 9 18 639,755 97,236 3,198.78 486.18 0.16 0.19 0.35 17 Chunk 39 107.75 127.5 147 268 46,978,904 2,079,562 234,894.52 10,397.81 11.74 4.16 15.90 17 Combined 3 6 28.5 127.25 268 47,618,659 2,176,798 119,046.65 5,442.00 11.90 4.35 16.26 19 Document 2 7 10 13 26 639,755 230,297 3,198.78 1,151.49 0.80 2.30 3.10 19 Chunk 25 115.75 131 150 209 19,159,387 3,141,812 95,796.94 15,709.06 23.95 31.42 55.37 19 Combined 2 10 25.5 130.5 209 19,799,142 3,372,109 49,497.86 8,430.27 24.75 33.72 58.47
@@
In Table 6, we report the full set of statistics for the runs selected in our feasibility analysis, covering all prompt variants and the ICL-5/TE3-S configuration within the non-agentic workflow. Additionally, we assess the practical feasibility of the proposed approach by analysing the cost distribution across the selected runs. Document ranking tasks consistently incur low total costs (below \$3), reflecting their modest inference requirements. In contrast, chunk ranking tasks dominate the overall expenditure, with total costs ranging from \$4.7 to \$55, depending on input size and reasoning depth. Combined task runs follow the same pattern, as their total costs are primarily driven by chunk ranking computations. Among all configurations, Runs 15 and 19 exhibit the highest total costs (\$57.9 and \$58.5 respectively) due to extensive token generation from the larger GPT-5 model, whereas earlier runs such as Runs 2 and 9 remain significantly more economical ( \$5) when using smaller models. These results confirm that document-level ranking remains cost efficient, while chunk reasoning is more expensive as it offers richer contextual understanding and remains the major contributor to overall cost.
A.6 Reproducibility Analysis
A.6.1 Descriptive Statistics

### v2

@@ A.10.1 Cost Analysis
Table 13: Latency, token usage, and cost statistics for different runs used in feasibility analysis on FinAgentBench.
@@
ID Task Latency Statistics (seconds) Token Usage Statistics Cost Statistics (USD) Min $Q_{1}$ Median $Q_{3}$ Max $\sum_{prompt}$ $\sum_{completion}$ $\mu_{prompt}$ $\mu_{completion}$ $\sum_{prompt}$ $\sum_{completion}$ $\sum_{combined}$ 2 Document 2 3 3 4 100 150,541 61,904 752.71 309.52 0.04 0.12 0.16 2 Chunk 4 12 32 43 4190 15,010,574 519,830 75,052.87 2,599.15 3.75 1.04 4.79 2 Combined 2 3 7 32 4190 15,161,115 581,734 37,902.79 1,454.34 3.79 1.16 4.95 9 Document 2 4 4 5 17 241,941 68,752 1,209.71 343.76 0.06 0.14 0.20 9 Chunk 6 13 32 47 99 14,963,931 490,556 74,819.66 2,452.78 3.74 0.98 4.72 9 Combined 2 4 8 32 99 15,205,872 559,308 38,014.68 1,398.27 3.80 1.12 4.92 12 Document 5 12 17 25 83 327,741 297,498 1,638.71 1,487.49 0.08 0.59 0.68 12 Chunk 9 62 136 194 665 15,214,202 2,271,311 76,071.01 11,356.56 3.80 4.54 8.35 12 Combined 5 17 38 135 665 15,541,943 2,568,809 38,854.86 6,422.02 3.89 5.14 9.02 13 Document 2 4 5 6 15 426,941 84,369 2,134.71 421.85 0.11 0.17 0.28 13 Chunk 34 92 105 117 171 19,200,124 1,734,012 96,000.62 8,670.06 4.80 3.47 8.27 13 Combined 2 5 25 105 171 19,627,065 1,818,381 49,067.66 4,545.95 4.91 3.64 8.54 15 Document 2 5 7 10 25 426,941 180,699 2,134.71 903.50 0.53 1.81 2.34 15 Chunk 23 133 156 179 240 19,170,218 3,159,981 95,851.09 15,799.91 23.96 31.60 55.56 15 Combined 2 7 24 156 240 19,597,159 3,340,680 48,992.90 8,351.70 24.50 33.41 57.90 17 Document 3 5 6 9 18 639,755 97,236 3,198.78 486.18 0.16 0.19 0.35 17 Chunk 39 108 128 147 268 46,978,904 2,079,562 234,894.52 10,397.81 11.74 4.16 15.90 17 Combined 3 6 29 127 268 47,618,659 2,176,798 119,046.65 5,442.00 11.90 4.35 16.26 19 Document 2 7 10 13 26 639,755 230,297 3,198.78 1,151.49 0.80 2.30 3.10 19 Chunk 25 116 131 150 209 19,159,387 3,141,812 95,796.94 15,709.06 23.95 31.42 55.37 19 Combined 2 10 26 131 209 19,799,142 3,372,109 49,497.86 8,430.27 24.75 33.72 58.47
@@
In Table 13, we report the full set of statistics for the runs selected in our feasibility analysis, covering all prompt variants and the ICL-5/TE3-S configuration within the non-agentic workflow. Additionally, we assess the practical feasibility of the proposed approach by analyzing the cost distribution across the selected runs. Document ranking tasks consistently incur low total costs (below \$3), reflecting their modest inference requirements. In contrast, chunk ranking tasks dominate the overall expenditure, with total costs ranging from \$4.7 to \$55, depending on input size and reasoning depth. Combined task runs follow the same pattern, as their total costs are primarily driven by chunk ranking computations. Among all configurations, Runs 15 and 19 exhibit the highest total costs (\$57.9 and \$58.5 respectively) due to extensive token generation from the larger GPT-5 model, whereas earlier runs such as Runs 2 and 9 remain significantly more economical ( \$5) when using smaller models. These results confirm that document-level ranking remains cost-efficient, while chunk reasoning is more expensive as it offers richer contextual understanding and remains the major contributor to overall cost.
Table 14: Comprehensive latency, token usage and cost statistics for different prompt versions and models on FiQA-2018.
Prompt Model ICL Total Prompt Total Completion Total Time (Ks) Time Per Query (s) Total Cost (USD) Tokens (M) Tokens (K) $P_{1}$ GPT-4.1 N/A 70.02 237.63 71.70 110.65 141.90 $P_{1}$ GPT-5-mini N/A 82.44 293.29 38.63 59.61 21.20 $P_{1}$ GPT-5 N/A 66.27 479.44 51.17 78.97 87.63 $P_{1}$ GPT-4.1 ICL-5 73.02 249.23 62.19 95.97 148.04 $P_{1}$ GPT-5-mini ICL-5 82.58 234.98 31.20 48.15 21.11 $P_{1}$ GPT-5 ICL-5 82.58 474.73 56.23 86.77 107.97 $P_{2}$ GPT-4.1 N/A 61.60 442.55 85.43 131.84 126.74 $P_{2}$ GPT-5-mini N/A 82.42 480.63 35.75 55.17 21.57 $P_{2}$ GPT-5 N/A 82.42 790.47 54.24 83.70 110.93 $P_{2}$ GPT-4.1 ICL-5 60.85 396.73 81.37 125.57 124.87 $P_{2}$ GPT-5-mini ICL-5 82.56 389.64 29.29 45.20 21.42 $P_{2}$ GPT-5 ICL-5 82.47 787.34 53.78 82.99 110.96 $P_{3}$ GPT-4.1 N/A 40.40 344.30 82.18 126.82 83.55 $P_{3}$ GPT-5-mini N/A 82.43 505.36 35.09 54.15 21.62 $P_{3}$ GPT-5 N/A 82.30 606.21 49.82 76.88 108.94 $P_{3}$ GPT-4.1 ICL-5 58.99 328.15 77.96 120.30 120.60 $P_{3}$ GPT-5-mini ICL-5 82.57 442.93 29.32 45.25 21.53 $P_{3}$ GPT-5 ICL-5 82.57 607.09 48.14 74.29 109.28 $P_{4}$ GPT-4.1 N/A 38.70 272.99 73.19 112.95 79.58 $P_{4}$ GPT-5-mini N/A 82.42 341.96 34.59 53.58 21.29 $P_{4}$ GPT-5 N/A 82.30 511.02 55.24 85.25 107.98 $P_{4}$ GPT-4.1 ICL-5 43.11 283.30 86.78 133.92 88.48 $P_{4}$ GPT-5-mini ICL-5 82.56 256.27 29.02 44.78 21.15 $P_{4}$ GPT-5 ICL-5 82.56 516.58 52.37 80.82 108.37
[…]
An interesting cost-performance paradox emerges when comparing GPT-5 to GPT-4.1. Despite being a more advanced model, GPT-5 consistently delivers superior performance (0.6130 vs. 0.4307 NDCG@10 for $P_{1}$ without ICL) while maintaining lower operational costs (\$88 vs. \$142). This counterintuitive result stems from GPT-5’s substantially more efficient token generation patterns. The key lies in GPT-5’s architectural improvements that enable more concise reasoning: although GPT-5 generates approximately twice as many completion tokens as GPT-4.1, its significantly lower per-token pricing more than compensates for this increase. Additionally, GPT-5’s faster processing time compared to GPT-4.1 suggests better computational efficiency at the infrastructure level. These findings challenge the common assumption that newer, more capable models necessarily incur higher costs, demonstrating instead that architectural efficiency gains can simultaneously improve both accuracy and cost-effectiveness. We do note that API pricing is a commercial decision that does not always track computational overhead, and these cost patterns may shift as pricing evolves. For organizations prioritizing maximum performance without cost constraints, GPT-5 represents the clear optimal choice, delivering state-of-the-art results at lower total cost than the previous generation model.
Table 15: Comprehensive latency, token usage and cost statistics for different prompt versions and models across various evaluation modes on FinanceBench.
Prompt Model Eval Mode ICL Total Prompt Total Completion Total Time Per Total Cost
Tokens (K) Tokens (K) Time (s) Query (s) (USD)
Baselines
Baseline GPT-4.1 Single Store N/A 4.81 27.13 494.94 3.30 0.23
Baseline GPT-4.1 Shared Store N/A 4.81 27.72 478.99 3.19 0.23
Baseline GPT-4.1 In Context N/A 1260.12 40.70 2466.77 16.45 25.53
Baseline GPT-4.1 Oracle N/A 159.80 27.08 329.78 2.20 0.54
Baseline GPT-5-mini Single Store N/A 4.81 18.39 1165.15 7.77 0.04
Baseline GPT-5-mini Shared Store N/A 4.81 18.69 1149.36 7.66 0.039
Baseline GPT-5-mini In Context N/A 1260.03 131.32 1505.37 10.04 3.41
Baseline GPT-5-mini Oracle N/A 158.90 138.23 1200.31 8.00 0.32
Baseline GPT-5 Single Store N/A 4.81 13.33 1556.93 10.38 0.14
Baseline GPT-5 Shared Store N/A 4.81 13.81 1674.66 11.16 0.14
Baseline GPT-5 In Context N/A 1260.03 147.64 2066.25 13.78 17.23
Baseline GPT-5 Oracle N/A 158.90 179.69 1433.36 9.56 2.00
PRISM
$P_{1}$ GPT-4.1 Single Store ICL-9 120.91 15.40 371.14 2.47 0.37
$P_{1}$ GPT-4.1 Shared Store ICL-9 154.02 14.59 332.71 2.22 0.42
$P_{1}$ GPT-4.1 In Context ICL-9 9885.98 36.66 2609.96 17.40 20.07
$P_{1}$ GPT-4.1 Oracle ICL-9 377.00 26.53 285.59 1.90 0.97
$P_{1}$ GPT-5-mini Single Store ICL-9 105.78 15.58 1649.79 11.00 0.06
$P_{1}$ GPT-5-mini Shared Store ICL-9 112.44 14.01 1579.32 10.53 0.06
$P_{1}$ GPT-5-mini In Context ICL-9 9868.73 130.71 1368.42 9.12 2.73
$P_{1}$ GPT-5-mini Oracle ICL-9 352.25 100.56 1104.98 7.37 0.29
$P_{1}$ GPT-5 Single Store ICL-9 138.01 13.68 1836.46 12.24 0.31
$P_{1}$ GPT-5 Shared Store ICL-9 134.41 12.32 1834.41 12.23 0.29
$P_{1}$ GPT-5 In Context ICL-9 9929.93 187.48 2586.63 17.24 14.29
$P_{1}$ GPT-5 Oracle ICL-9 341.00 166.67 1683.26 11.22 2.09
$P_{1}$ GPT-4.1 Single Store N/A 4.81 17.99 404.56 2.70 0.15
$P_{1}$ GPT-4.1 Shared Store N/A 4.81 16.73 373.95 2.49 0.14
$P_{1}$ GPT-4.1 In Context N/A 9744.98 39.30 1584.33 10.56 19.80
$P_{1}$ GPT-4.1 Oracle N/A 192.80 26.96 281.26 1.88 0.60
$P_{1}$ GPT-5-mini Single Store N/A 4.81 17.12 1660.00 11.07 0.04
$P_{1}$ GPT-5-mini Shared Store N/A 4.81 16.22 1574.42 10.50 0.03
$P_{1}$ GPT-5-mini In Context N/A 9744.53 147.37 1484.62 9.90 2.73
$P_{1}$ GPT-5-mini Oracle N/A 192.35 113.48 1274.31 8.50 0.28
$P_{1}$ GPT-5 Single Store N/A 4.81 14.16 1872.66 12.48 0.15
$P_{1}$ GPT-5 Shared Store N/A 4.81 13.98 1914.46 12.76 0.15
$P_{1}$ GPT-5 In Context N/A 9744.53 191.30 2651.92 17.68 14.09
$P_{1}$ GPT-5 Oracle N/A 192.35 187.49 1887.10 12.58 2.12
$P_{2}$ GPT-4.1 Single Store ICL-9 124.14 22.52 501.14 3.34 0.43
$P_{2}$ GPT-4.1 Shared Store ICL-9 104.34 18.15 351.93 2.35 0.35
$P_{2}$ GPT-4.1 In Context ICL-9 9859.73 41.07 1492.38 9.95 20.05
$P_{2}$ GPT-4.1 Oracle ICL-9 335.20 28.47 468.11 3.12 0.90
$P_{2}$ GPT-5-mini Single Store ICL-9 124.14 20.94 1636.65 10.91 0.07
$P_{2}$ GPT-5-mini Shared Store ICL-9 127.56 19.75 1548.20 10.32 0.07
$P_{2}$ GPT-5-mini In Context ICL-9 9891.23 157.62 1600.95 10.67 2.79
$P_{2}$ GPT-5-mini Oracle ICL-9 343.40 125.91 1496.93 9.98 0.34
$P_{2}$ GPT-5 Single Store ICL-9 132.06 18.57 1867.68 12.45 0.35
$P_{2}$ GPT-5 Shared Store ICL-9 132.96 17.82 1937.48 12.92 0.34
$P_{2}$ GPT-5 In Context ICL-9 9912.68 186.49 1710.18 11.40 14.26
$P_{2}$ GPT-5 Oracle ICL-9 378.80 176.42 1251.60 8.34 2.24
$P_{2}$ GPT-4.1 Single Store N/A 4.81 21.43 433.74 2.89 0.18
$P_{2}$ GPT-4.1 Shared Store N/A 4.81 20.33 402.21 2.68 0.17
$P_{2}$ GPT-4.1 In Context N/A 9742.13 44.08 1469.06 9.79 19.84
$P_{2}$ GPT-4.1 Oracle N/A 189.95 30.35 335.19 2.23 0.62
$P_{2}$ GPT-5-mini Single Store N/A 4.81 23.49 1678.87 11.19 0.05
$P_{2}$ GPT-5-mini Shared Store N/A 4.81 23.33 1723.72 11.49 0.05
$P_{2}$ GPT-5-mini In Context N/A 9741.68 171.78 1622.31 10.82 2.78
$P_{2}$ GPT-5-mini Oracle N/A 189.50 143.25 1590.80 10.61 0.33
$P_{2}$ GPT-5 Single Store N/A 4.81 19.92 2076.89 13.85 0.21
$P_{2}$ GPT-5 Shared Store N/A 4.81 19.27 2047.74 13.65 0.20
$P_{2}$ GPT-5 In Context N/A 9741.68 198.25 1794.61 11.96 14.16
$P_{2}$ GPT-5 Oracle N/A 189.50 191.22 1325.28 8.84 2.15
$P_{3}$ GPT-4.1 Single Store ICL-9 90.31 19.62 379.90 2.53 0.34
$P_{3}$ GPT-4.1 Shared Store ICL-9 128.82 23.29 430.37 2.87 0.44
$P_{3}$ GPT-4.1 In Context ICL-9 9908.78 55.20 1596.56 10.64 20.26
$P_{3}$ GPT-4.1 Oracle ICL-9 337.70 38.70 478.79 3.19 0.99
$P_{3}$ GPT-5-mini Single Store ICL-9 128.82 22.09 1444.10 9.63 0.08
$P_{3}$ GPT-5-mini Shared Store ICL-9 115.14 21.04 1457.49 9.72 0.07
$P_{3}$ GPT-5-mini In Context ICL-9 9896.18 184.37 1852.82 12.35 2.84
$P_{3}$ GPT-5-mini Oracle ICL-9 330.35 134.41 1502.07 10.01 0.35
$P_{3}$ GPT-5 Single Store ICL-9 130.81 18.69 2000.21 13.33 0.35
$P_{3}$ GPT-5 Shared Store ICL-9 139.81 18.96 2061.80 13.75 0.36
$P_{3}$ GPT-5 In Context ICL-9 9882.83 200.39 1768.46 11.79 14.36
$P_{3}$ GPT-5 Oracle ICL-9 354.50 187.28 1244.50 8.30 2.32
$P_{3}$ GPT-4.1 Single Store N/A 4.81 24.26 466.67 3.11 0.20
$P_{3}$ GPT-4.1 Shared Store N/A 4.81 22.01 423.23 2.82 0.19
$P_{3}$ GPT-4.1 In Context N/A 9744.83 59.43 2204.91 14.70 19.97
$P_{3}$ GPT-4.1 Oracle N/A 192.65 40.32 491.83 3.28 0.71
$P_{3}$ GPT-5-mini Single Store N/A 4.81 26.19 1777.55 11.85 0.05
$P_{3}$ GPT-5-mini Shared Store N/A 4.81 24.58 1448.53 9.66 0.05
$P_{3}$ GPT-5-mini In Context N/A 9744.38 195.59 1901.49 12.68 2.83
$P_{3}$ GPT-5-mini Oracle N/A 192.20 162.52 1675.36 11.17 0.37
$P_{3}$ GPT-5 Single Store N/A 4.81 21.22 2072.32 13.82 0.22
$P_{3}$ GPT-5 Shared Store N/A 4.81 20.98 2210.27 14.74 0.22
$P_{3}$ GPT-5 In Context N/A 9744.38 213.21 1876.14 12.51 14.31
$P_{3}$ GPT-5 Oracle N/A 192.20 222.01 1449.19 9.66 2.46
$P_{4}$ GPT-4.1 Single Store ICL-9 111.54 27.35 403.09 2.69 0.44
$P_{4}$ GPT-4.1 Shared Store ICL-9 115.51 28.06 390.28 2.60 0.46
$P_{4}$ GPT-4.1 In Context ICL-9 9908.48 79.75 1720.44 11.47 20.45
$P_{4}$ GPT-4.1 Oracle ICL-9 333.65 50.99 535.05 3.57 1.08
$P_{4}$ GPT-5-mini Single Store ICL-9 127.21 44.54 1950.23 13.00 0.12
$P_{4}$ GPT-5-mini Shared Store ICL-9 124.86 43.19 2001.19 13.34 0.12
$P_{4}$ GPT-5-mini In Context ICL-9 9907.43 240.14 2291.46 15.28 2.96
$P_{4}$ GPT-5-mini Oracle ICL-9 319.55 176.88 1747.33 11.65 0.43
$P_{4}$ GPT-5 Single Store ICL-9 118.21 24.77 2019.59 13.46 0.40
$P_{4}$ GPT-5 Shared Store ICL-9 136.02 23.82 2165.59 14.44 0.41
$P_{4}$ GPT-5 In Context ICL-9 9905.78 231.55 2216.82 14.78 14.70
$P_{4}$ GPT-5 Oracle ICL-9 327.20 234.87 1777.46 11.85 2.76
$P_{4}$ GPT-4.1 Single Store N/A 4.81 28.82 521.27 3.48 0.24
$P_{4}$ GPT-4.1 Shared Store N/A 4.81 27.21 484.88 3.23 0.23
$P_{4}$ GPT-4.1 In Context N/A 9746.48 82.14 1714.16 11.43 20.15
$P_{4}$ GPT-4.1 Oracle N/A 194.30 51.59 522.74 3.48 0.80
$P_{4}$ GPT-5-mini Single Store N/A 4.81 42.44 2156.44 14.38 0.09
$P_{4}$ GPT-5-mini Shared Store N/A 4.81 47.24 2128.26 14.19 0.10
$P_{4}$ GPT-5-mini In Context N/A 9746.03 275.14 2597.60 17.32 2.99
$P_{4}$ GPT-5-mini Oracle N/A 193.85 194.71 2007.18 13.38 0.44
$P_{4}$ GPT-5 Single Store N/A 4.81 28.72 2761.10 18.41 0.29
$P_{4}$ GPT-5 Shared Store N/A 4.81 31.62 2893.80 19.29 0.32
$P_{4}$ GPT-5 In Context N/A 9746.03 246.76 2276.21 15.17 14.65
$P_{4}$ GPT-5 Oracle N/A 193.85 263.13 2043.63 13.62 2.87
Table LABEL:tab:full_financebench_analysis presents comprehensive computational cost analysis across all prompt variants, models, and evaluation modes on FinanceBench dataset. Single Store and Shared Store modes consistently incur minimal costs that are below \$0.50 for all models, reflecting their limited context requirements with prompt tokens. In contrast, In Context mode dominates the overall expenditure, with costs ranging from \$2.73 to \$25.53 depending on the model, primarily driven by the extensive prompt token requirements needed to embed full document context within the prompt. Oracle mode occupies a middle ground with costs between \$0.28 and \$2.87, as it requires moderate context to provide relevant passages without full document inclusion.
Model selection significantly impacts cost-performance trade-offs. GPT-5-mini emerges as the most cost-efficient option across all evaluation modes, with Oracle mode costing merely \$0.28–\$0.44 compared to \$0.60–\$2.87 for GPT-5 and \$0.80–\$1.08 for GPT-4.1 under similar conditions. Despite this dramatic cost advantage, GPT-5-mini maintains superior performance, achieving 98% accuracy in several configurations. GPT-5, while incurring higher costs, consistently outperforms other models in most configurations. GPT-4.1 exhibits the least favorable cost-performance profile, with comparable costs to GPT-5 but consistently lower accuracy.
Prompt complexity introduces additional computational overhead with varying returns. The straightforward prompt $P_{1}$ achieves optimal cost-efficiency, maintaining low completion token generation while delivering top-tier performance. More complex prompts like $P_{4}$ generate substantially more completion tokens, increasing costs without proportional performance gains in most evaluation modes. The impact of ICL is particularly pronounced in In Context mode, where it reduces costs slightly for GPT-5 by providing better task guidance that reduces completion token generation.
Processing time analysis reveals GPT-4.1 as the fastest model in Single Store and Shared Store modes but shows significant variability in other modes. GPT-5 and GPT-5-mini show consistent processing times across different modes, with GPT-5-mini being slightly faster. These results confirm that for production deployment on QA tasks, In Context mode’s substantial cost premium makes it impractical for large-scale applications despite marginal performance benefits. GPT-5-mini with simpler prompts offers the optimal balance of accuracy, cost, and latency.
A.11 Reproducibility Analysis
As reported in Tables 7 and 7, all runs exhibit low variability with standard deviations ($s<0.011$) and coefficients of variation (CV $1.6\%$), indicating stable and reproducible outcomes. The narrow 95% confidence intervals (CI) confirm the statistical reliability of the mean performance estimates. Run 12 recorded the lowest mean score (0.68005), while Runs 15–19 achieved consistently higher averages, with non-overlapping CIs showing statistically significant gains following configuration updates. Welch’s two-sample $t$-tests further support these improvements: Run 12 vs. Run 15 (t=3.969t=-3.969, $p=0.0057$), Run 12 vs. Run 18 (t=4.981t=-4.981, $p=0.0022$), and Run 12 vs. Run 19 (t=6.582t=-6.582, $p=0.0014$). The progressively decreasing $p$-values and increasing $t$-statistics indicate that each subsequent configuration yields more pronounced improvements. In general, the low variance and narrow confidence bounds demonstrate that the updated configurations produce statistically significant and reproducible results.
A.11.1 Descriptive Statistics